# Language Model Agents Suffer from Compositional Generalization in Web Automation

### v1

@@ 6.3 Do Advanced LLMs Solve Compositional Tasks?
Figure 4 presents the results when we adopt other advanced LLMs, than gpt-3.5-turbo, as a backbone of each LMA. The more capable models, such as gpt-4 in a generalist aspect and text-davinci-003 in a code generation, can improve the success rate of all the prompted LMAs. However, even gpt-4 is still far from the generalization in compositional tasks (from 28.7% to 56.0% by RCI) or from dealing with reverse-order instructions (from 19.2% to 43.5% by RCI). This indicates that we need much better foundation models to realize deployable LMA in complex real-world decision-making tasks. We provide failure examples in Appendix G.
6.4 What Determines Task Complexity in Web Automation?
Figure 5 visualizes the correlation between the success rate averaged across WebGUM, HTML-T5, RCI, AdaPlanner, and Synapse (y-axis) and each statistic of compositional tasks (x-axis), such as synthesized success rate – a product of base task success rates among compositional tasks – the number of instruction tokens, and max depth of HTML subtrees. Synthesized success rate positively correlates with an average success rate ($R=0.691$), indicating that compositional task difficulty takes over base task difficulties. In addition, the number of instruction tokens ($R=-0.579$) and the max depth of HTML subtrees ($R=-0.433$) show negative correlations. All those are statistically significant in paired t-test with $p<0.01$. In contrast, other task statistics, such as synthesized success rate with human performance, the number of HTML tokens, and elements in HTML, just show relatively weaker correlations (see Appendix F for the details). This analysis suggests that HTML with larger depth and long instructions make generalizing compositional tasks challenging. The complexity of HTML is determined by its depth rather than its length or the number of elements. This might come from the hierarchical nature of HTML; in deeper HTML subtrees, the elements near the root tend to be distant from each other after the traversal. Such sparsity may cause confusion during planning.
@@ 7 Discussion
Generalizable Prompting Methods  The results of Synapse and RCI in Figure 2 imply that those prompted LMAs have some “over-fitting” trends to the base MiniWoB tasks. While the robustness across the prompts has been investigated in natural language tasks (Wei et al., 2022; Kojima et al., 2022), it is not well understood in the decision making problems. Because we will not be able to prepare the optimal self-improvement iterations or decomposed prompts for all the possible instructions and task compositions, even if using optimal exemplar retrievers, we should care more about the generalization of prompting methods for the agent systems.
Agent-Specialized Large Language Models  As shown in Figure 4, the more capable LLMs, such as gpt-4, can improve the performance of LMAs in CompWoB. However, it has not reached the base MiniWoB yet (e.g. from 90.6% to 56.0% in RCI, and from 98.5% to 41.4% in Synapse). Similarly, as described in Section 4.4, transferred LMAs can perform better if the training dataset has a good balance and coverage, but it is far from sufficient compositional generalization or instruction generalization. The current pre-trained LLMs may still not be sufficient to generalize to complex decision making tasks, and then, in addition to prompting methods, the development of agent-specialized LLMs with enhanced reasoning and generalization would be expected.
@@ Acknowledgments
We thank Mustafa Safdari, Yusuke Iwasawa, Heiga Zen, and Doina Precup for helpful feedback on this work. HF was supported by JSPS KAKENHI Grant Number JP22J21582.
@@ Appendix E Details of MiniWoB Dataset
To resolve the data-imbalance problem, we first run Synapse (Zheng et al., 2023) on MiniWoB and collect 77K additional demonstrations across 16 tasks on top of 347K demonstrations (Furuta et al., 2023) to compensate for the lack of data in specific tasks (Strategy A). We use PaLM 2-L (Anil et al., 2023) as a backbone LLM for Synapse. We then reduce the number of demonstrations for the tasks the agents can solve to focus on more challenging tasks. Based on brute-force task complexity (Appendix D), we consider the following four strategies:
@@
Through the empirical evaluations (Section 4.4), we find that Strategy D realizes a well-balanced dataset to improve the performance.
@@ Appendix G Failure Examples with Advanced Language Models
We provide several failure episodes with advanced language models such as gpt-4 and text-davinci-003 in Table 11. The left columns have correct plans, and the right columns have failure plans. Compared to gpt-3.5-turbo (Table 2), LMAs based on advanced models may not ignore the necessary intermediate actions. However, they tend to suffer from tiny errors such as capitalization (RCI, AdaPlanner) or attributes in HTML (Synapse).
@@ Appendix I Prompted Language Model Agents with Oracle Exemplars
We here evaluate the performance of RCI with gpt-4 and oracle exemplars on 20 two-way tasks in CompWoB. We provide two demonstrations per task in the prompts. Figure 7 shows that RCI with gpt-4 and oracle exemplars achieves 82.9% success rate, which is the best among the baselines, such as HTML-T5++ (73.9%), RCI with combination exemplars (gpt-3.5-turbo: 46.9%, gpt-4: 71.5%). See Table 13 for other baselines. This ensures that if a prompt includes how to perform on compositional tasks, the performance gets better. However, please also keep in mind that providing exemplars for every compositional problem is infeasible given the huge space of problems and the cost of collecting examples for every one of them.
Figure 7: Average success rate of LMAs in 20 two-way tasks from CompWoB. RCI with gpt-4 achieves the best performance when the oracle exemplars are provided (82.9%) in the prompt.

### v3

@@ 6.3 Do Advanced LLMs Solve Compositional Tasks?
Figure 4 presents the results when we adopt other advanced LLMs, than gpt-3.5-turbo, as a backbone of each LMA. The more capable models, such as gpt-4 in a generalist aspect and text-davinci-003 in a code generation, can improve the success rate of all the prompted LMAs. However, even gpt-4 is still far from the generalization in compositional tasks (from 28.7% to 56.0% by RCI) or from dealing with reverse-order instructions (from 19.2% to 43.5% by RCI). This indicates that we need much better LLMs to realize deployable systems in complex real-world decision-making tasks. We provide failure examples in Appendix G.
6.4 What Determines Task Complexity on the Web?
Figure 5 visualizes the correlation between the success rate averaged across WebGUM, HTML-T5, RCI, AdaPlanner, and Synapse (y-axis) and each statistic of compositional tasks (x-axis), such as synthesized success rate – a product of base task success rates among compositional tasks – the number of instruction tokens, and max depth of HTML subtrees. Synthesized success rate positively correlates with an average success rate ($R=0.691$), indicating that compositional task difficulty takes over base task difficulties. In addition, the number of instruction tokens ($R=-0.579$) and the max depth of HTML subtrees ($R=-0.433$) show negative correlations. All those are statistically significant in paired t-test with $p<0.01$. In contrast, other task statistics, such as synthesized success rate with human performance, the number of HTML tokens, and elements in HTML, just show relatively weaker correlations (see Appendix F for the details). This analysis suggests that HTML with larger depth and long instructions make generalizing compositional tasks challenging. The complexity of HTML is determined by its depth rather than its length or the number of elements. This might come from the hierarchical nature of HTML; in deeper HTML subtrees, the elements near the root tend to be distant from each other after the traversal. Such sparsity may cause confusion during planning. We also report the individual characteristics of each agent in Appendix J.
@@ 7 Discussion
Generalizable Prompting Methods  The results of Synapse and RCI in Figure 2 imply that those prompted LMAs have “over-fitting” trends to the base MiniWoB tasks. While the robustness across the prompts has been investigated in natural language tasks (Wei et al., 2022; Kojima et al., 2022), it is not well understood in decision making. Because we will not be able to prepare the optimal self-improvement iterations or decomposed prompts for all the possible instructions and task compositions, even if using optimal exemplar retrievers, we should care more about the generalization of prompting methods for the agent systems.
Agent-Specialized Large Language Models  As shown in Figure 4, the more capable LLMs, such as gpt-4, can improve the performance of LMAs in CompWoB. However, it has not reached the base MiniWoB yet (e.g. from 90.6% to 56.0% in RCI, and from 98.5% to 41.4% in Synapse). Similarly, as described in Section 4.4, transferred LMAs can perform better if the training dataset has a good balance and coverage, but it is far from sufficient generalization to compositionality and instructions. The current pre-trained LLMs may still not be sufficient to generalize to complex decision making tasks, and then, in addition to prompting methods, the development of agent-specialized LLMs with enhanced reasoning and generalization through the instruction-tuning (Zeng et al., 2023) or RLHF/AIF (Furuta et al., 2024) would be expected.
@@ 8 Conclusion
Ethics Statements
This paper systematically evaluates the performance of existing language model agents for web automation in realistic compositional tasks, and unveils remaining challenges towards broader generalization and real-world deployments. This technique could realize capable AI assistant tools on digital devices (e.g. computers or smartphones), and improve productivity and accessibility for society.
While we anticipate the positive aspects of autonomous agents, for responsible development, we should also consider the potential harmful applications and unintended consequences. The misuse of web automation would threaten cyber security, and the users may get scammed. To reduce these risks, it is essential for researchers, policymakers, and industry to discuss concrete guidelines and regulations.
@@ Acknowledgments
We thank Mustafa Safdari, Yingjie Miao, Yusuke Iwasawa, Heiga Zen, and Doina Precup for the help on this work. HF was supported by JSPS KAKENHI Grant Number JP22J21582.
@@ Appendix B Pseudo Code for Prompted Language Model Agents
We explicitly distinguish LLM itself and LMA (combination of LLM, prompting and pipeline) in this paper. We provide the pseudo code for prompted LMAs to clarify their methodological difference. RCI (Kim et al., 2023) is the first to use prompting in a self-refinement loop, outperforming imitation- or reinforcement-learned agents on MiniWoB benchmark that requires millions of demonstrations to work. AdaPlanner (Sun et al., 2023) and Synapse (Zheng et al., 2023) are the follow-up works outperforming RCI via code generation from environmental feedback or via well-designed decomposed prompts with retrieval.
@@ Appendix E Details of MiniWoB Dataset
To resolve the data-imbalance problem, we first run Synapse (Zheng et al., 2023) on MiniWoB and collect 77K additional demonstrations across 16 tasks on top of 347K demonstrations (Furuta et al., 2023) to compensate for the lack of data in specific tasks (Strategy A). We use PaLM 2-L (Anil et al., 2023) as a backbone LLM for Synapse. We then reduce the number of demonstrations for the tasks the agents can solve to focus on more challenging tasks. Based on brute-force task complexity (Appendix D), we consider the following four strategies as shown in Table 11:
@@
We hold out 21K episodes (5% of 347K + 77K = 424K) as a validation split, and after the convergence, we choose the top-5 checkpoints that achieve higher offline validation accuracy, run those checkpoints online on MiniWoB benchmarks, and then report the best success rate. Through the empirical evaluations (Section 4.4), we find that Strategy D realizes a well-balanced dataset to improve the performance. Except for these dataset mixtures, we follow Gur et al. (2023) for other training hyper-parameters of HTML-T5++. We have used cloud TPU-v3, which has a 32 GiB HBM memory space, with 128 cores. Each finetuning experiment takes about 1-2 days.
@@ Appendix G Failure Examples with Advanced Language Models
We provide several failure episodes with advanced language models such as gpt-4 and text-davinci-003 in Table 12. The left columns have correct plans, and the right columns have failure plans. Qualitatively, LMAs based on advanced models manage to reduce the ratio of failure modes in gpt-3.5-turbo (Table 3), such as skipping necessary intermediate steps and hallucination of unnecessary actions. However, they tend to suffer from tiny errors more such as capitalization (RCI, AdaPlanner) or attributes in HTML (Synapse).
@@ Appendix H Per-Task Performance on MiniWoB and CompWoB
Task RCI (optimal) Synapse (optimal) AdaPlanner (optimal) RCI (oracle, gpt-4) RCI (zero) RCI (first) RCI (second) RCI (comb) RCI (gpt-4) Synapse (first) Synapse (second) Synapse (best) Synapse (f;gpt-4) Synapse (s;gpt-4) Synapse (b;gpt-4) AdaPlanner AdaPlanner (davinci)
click-button_click-checkboxes 1.000 1.000 1.000 1.000 0.370 0.710 0.730 0.800 0.880 0.810 0.840 0.840 0.980 0.840 0.980 0.060 0.890
click-button_click-checkboxes-transfer 1.000 1.000 1.000 0.980 0.270 0.790 0.790 0.810 1.000 0.540 0.740 0.740 0.960 0.740 0.960 0.000 0.940
click-button_click-dialog 1.000 1.000 1.000 1.000 0.740 0.830 0.820 0.940 1.000 1.000 0.650 1.000 1.000 0.660 1.000 0.900 0.980
click-button_click-link 1.000 1.00 0.980 0.960 0.870 0.830 0.870 0.920 0.920 0.990 0.000 0.990 1.000 0.000 1.000 0.940 0.970
click-button_click-option 1.000 1.000 1.000 1.000 0.580 0.650 0.370 0.710 0.420 0.920 0.870 0.920 0.960 0.880 0.960 0.000 0.860
click-button-sequence_click-checkboxes 1.000 1.000 1.000 0.980 0.440 0.520 0.820 0.670 0.820 0.710 0.890 0.890 0.940 0.840 0.940 0.060 0.680
click-button-sequence_click-option 1.000 1.000 1.000 1.000 0.540 0.830 0.620 0.760 0.800 0.900 0.000 0.900 0.900 0.000 0.900 0.000 0.490
click-button-sequence_login-user-popup 1.000 1.000 1.000 1.000 0.000 0.000 0.000 0.000 0.100 0.000 0.000 0.000 0.300 0.000 0.300 0.420 0.000
click-link_click-button 1.000 1.000 0.980 0.980 0.790 0.830 0.970 0.980 0.980 0.010 1.000 1.000 0.000 1.000 1.000 0.950 0.800
click-link_click-dialog 1.000 1.000 0.980 0.660 0.200 0.200 0.460 0.600 0.640 0.000 0.000 0.000 0.000 0.000 0.000 0.540 0.490
click-link_click-widget 0.980 1.000 0.980 1.000 0.550 0.520 0.820 0.900 0.780 0.370 0.000 0.370 0.820 0.000 0.820 0.980 0.890
click-link_enter-text 1.000 1.000 0.960 1.000 0.090 0.160 0.200 0.060 0.840 0.000 0.000 0.000 0.000 0.000 0.000 0.860 0.920
click-option_enter-text 1.000 1.000 0.980 1.000 0.380 0.130 0.410 0.500 0.880 0.000 0.000 0.000 1.000 0.000 1.000 0.770 0.960
click-option_login-user 1.000 1.000 1.000 1.000 0.010 0.000 0.000 0.000 0.860 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
click-option_navigate-tree 0.860 0.960 0.820 0.960 0.550 0.470 0.500 0.720 0.820 0.000 0.240 0.240 0.560 0.960 0.960 0.620 0.590
click-widget_enter-password 0.980 1.000 0.980 0.760 0.000 0.000 0.000 0.000 0.740 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
click-widget_multi-layouts 0.706 0.940 0.840 0.400 0.000 0.000 0.000 0.000 0.240 0.000 0.760 0.760 0.000 1.000 1.000 0.000 0.000
enter-password_click-option 1.000 1.000 0.980 0.980 0.000 0.020 0.000 0.000 0.780 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
login-user_navigate-tree 0.860 0.960 0.820 0.780 0.020 0.010 0.000 0.000 0.700 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
multi-layouts_login-user 0.720 0.940 0.840 0.100 0.000 0.000 0.000 0.000 0.100 0.720 0.000 0.720 1.000 0.000 1.000 0.010 0.000
Average (two-way) 0.939 0.990 0.955 0.833 0.320 0.375 0.419 0.469 0.715 0.349 0.300 0.469 0.521 0.346 0.641 0.356 0.523
click-button_click-option_login-user 1.000 1.000 1.000 0.000 0.340 0.000 0.000 0.120 0.000 0.000 0.000 0.940 0.280 0.940 0.140 0.000
click-button-sequence_click-option_login-user 1.000 1.000 1.000 0.520 0.740 0.680 0.840 0.560 0.720 0.000 0.720 0.660 0.000 0.660 0.000 0.000
click-checkboxes_click-widget_click-button-sequence 0.980 1.000 1.000 0.400 0.660 0.660 0.820 1.000 0.820 0.000 0.820 0.820 0.180 0.820 0.580 0.720
click-checkboxes-transfer_click-button-sequence_enter-password 1.000 1.000 0.960 0.000 0.000 0.000 0.000 0.660 0.000 0.000 0.000 0.000 0.500 0.500 0.000 0.000
click-checkboxes-transfer_enter-password_click-dialog 1.000 1.000 0.960 0.000 0.000 0.000 0.000 0.280 0.000 0.420 0.420 0.000 0.400 0.400 0.000 0.000
click-dialog_click-button-sequence_enter-password 1.000 1.000 0.980 0.000 0.000 0.000 0.000 0.700 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
click-dialog_click-checkboxes-transfer_click-widget 0.980 1.000 0.980 0.140 0.540 0.340 0.400 0.680 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
click-link_click-button_click-dialog 1.000 1.000 0.980 0.240 0.270 0.480 0.610 0.600 0.000 0.600 0.600 0.000 0.620 0.620 0.560 0.580
click-widget_click-option_click-dialog 0.980 1.000 1.000 0.120 0.000 0.020 0.300 0.400 0.000 0.000 0.000 0.020 0.000 0.020 0.000 0.000
enter-password_click-checkboxes_login-user-popup 0.680 1.000 0.960 0.000 0.000 0.000 0.000 0.700 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
Average (three-way) 0.962 1.000 0.982 0.142 0.255 0.218 0.297 0.570 0.154 0.102 0.256 0.244 0.198 0.396 0.128 0.130
click-button-sequence_click-widget_click-link_click-button_click-checkboxes_click-option_click-dialog 0.980 1.000 0.980 0.160 0.020 0.380 0.080 0.720 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
click-button-sequence_click-widget_click-link_click-button_click-checkboxes_click-option_click-dialog_login-user 0.980 1.000 0.980 0.000 0.000 0.000 0.000 0.620 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
click-link_click-button_click-checkboxes_click-dialog 1.000 1.000 0.980 0.150 0.230 0.320 0.370 0.540 0.000 0.310 0.310 0.000 0.600 0.600 0.050 0.600
click-link_click-button_click-checkboxes_click-option_click-dialog 1.000 1.000 0.980 0.200 0.180 0.370 0.360 0.560 0.000 0.060 0.060 0.000 0.620 0.620 0.020 0.490
click-widget_click-link_click-button_click-checkboxes_click-option_click-dialog 0.980 1.000 0.980 0.140 0.360 0.100 0.300 0.480 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
Average (n-way) 0.988 1.000 0.980 0.130 0.158 0.234 0.222 0.584 0.000 0.074 0.074 0.000 0.244 0.244 0.014 0.218
click-checkboxes-transfer_multi-layouts_email-inbox-forward-nl-transition 0.720 0.940 0.823 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.220 0.440
click-option_login-user-transition 1.000 1.000 1.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
click-option_multi-layouts_click-widget_login-user-popup-transition 0.480 0.940 0.823 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
login-user_navigate-tree-transition 0.860 0.960 0.820 0.000 0.000 0.000 0.000 0.460 0.000 0.000 0.000 0.340 0.000 0.340 0.000 0.000
login-user-popup_email-inbox-forward-nl-turk-transition 0.639 1.000 0.980 0.000 0.000 0.200 0.000 0.120 0.000 0.000 0.000 0.000 0.000 0.000 0.700 0.560
Average (transition) 0.740 0.968 0.889 0.000 0.000 0.040 0.000 0.116 0.000 0.000 0.000 0.068 0.000 0.068 0.184 0.200
click-button_click-tab-2-hard 0.760 0.960 0.780 0.080 0.120 0.260 0.340 0.580 0.280 0.040 0.280 0.960 0.320 0.960 0.260 0.280
click-button-sequence_use-autocomplete 0.580 0.980 0.880 0.200 0.060 0.640 0.160 0.820 0.000 0.000 0.000 0.420 0.000 0.420 0.000 0.000
click-checkboxes-soft_enter-password 0.720 1.000 0.784 0.000 0.000 0.000 0.000 0.820 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
click-checkboxes-soft_multi-layouts 0.518 0.940 0.672 0.000 0.000 0.000 0.000 0.120 0.000 0.300 0.000 0.000 0.820 0.820 0.000 0.000
click-dialog_search-engine 1.000 1.000 1.000 0.100 0.020 0.000 0.000 0.920 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
click-dialog-2_click-widget 0.980 1.000 1.000 0.000 0.000 0.000 0.040 0.020 0.100 0.000 0.100 0.180 0.000 0.180 0.160 0.100
click-dialog-2_login-user-popup 0.680 1.000 0.980 0.000 0.000 0.000 0.000 0.060 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
click-widget_click-checkboxes-soft 0.706 1.000 0.800 0.080 0.100 0.060 0.200 0.600 0.000 0.000 0.000 0.000 0.000 0.000 0.100 0.140
enter-date_login-user 0.960 1.000 1.000 0.000 0.020 0.000 0.060 0.500 0.000 0.000 0.000 0.000 0.000 0.000 0.380 0.380
use-autocomplete_click-dialog 0.580 0.980 0.880 0.000 0.080 0.000 0.080 0.060 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
Average (two-way, easy-medium) 0.748 0.986 0.878 0.046 0.040 0.096 0.088 0.450 0.038 0.034 0.038 0.156 0.114 0.238 0.090 0.090
Average (total) 0.891 0.990 0.941 0.179 0.225 0.258 0.287 0.560 0.178 0.154 0.254 0.295 0.225 0.414 0.206 0.295
Table 16: Optimal success rate on CompWoB (synthetically calculated as a product of base task success rate in MiniWoB).
@@ Appendix I Prompted Language Model Agents with Oracle Exemplars
In this paper, we assume, as a realistic and important constraint, that: (1) prompted or finetuned LMAs are deployed as a service in the real world; (2) users are only allowed to interact with the agents via providing task instruction, and (3) not allowed to intervene the backend system or prompts. Providing exemplars or preparing finetuning demonstrations for every compositional problem is infeasible given the huge space of web automation problems. Moreover, it significantly hinders the user experience that the agents would be good at some specific format of instructions and that their violation of those affects the performance even if semantically the same.
As a sanity check of the environments and baseline agents, we here demonstrate that prompted LMAs can change their behaviors adaptively by modifying their prompts and demonstrations. We evaluate the performance of RCI with gpt-4 and oracle exemplars on 20 two-way tasks in CompWoB. We provide two demonstrations per task in the prompts. Figure 7 shows that RCI with gpt-4 and oracle exemplars achieves 82.9% success rate, which is the best among the baselines, such as HTML-T5++ (73.9%), RCI with gpt-3.5-turbo and oracle exemplars (56.8%), RCI with combination exemplars (gpt-3.5-turbo: 46.9%, gpt-4: 71.5%). See Table 14 for other baselines. This ensures that the tasks are feasible, and that if a prompt includes how to perform on compositional tasks, the performance gets better. In contrast, while oracle demonstrations help improve performance, they could not fully resolve the issues from reverse-order instructions.
Through the experiments in Section 6, we have intended to shed light on their zero-shot generalization capabilities of dealing with unknown sequential-task compositions. For reliable and robust web automation agents, we might need to leverage both prompting and finetuning approaches at the different development stages. For instance, prompted LMAs might work well as adaptive automated data collectors to the novel domains and finetuned LMAs as deployed agents for service.
Figure 7: Average success rate of LMAs in 20 two-way tasks from CompWoB. RCI with gpt-4 achieves the best performance when the oracle exemplars are provided (82.9%) in the prompt. While oracle demonstrations help improve performance, they could not fully resolve the issues from reverse-order instructions (for instance, 82.9% $\rightarrow$ 71.8% in RCI with gpt-4).
@@
Following the recent work in deep reinforcement learning literature (Furuta et al., 2021), we measure average performance as a proxy of oracle task solvability. If many kinds of agents perform poorly, such tasks are regarded as challenging. This can shed light on“task” or “environment” themselves, rather than “language model agents” or “prompting methods”, while such an analysis has been overlooked so far. Additionally, while the trend with average might be the same, distributional characteristics for each language model agent could be different. We extend our analysis by reporting the individual performances of each agent in Figure 8. Our results still indicate that while all the language model agents (HTML-T5++, RCI, AdaPlanner, Synapse) often show negative correlations between the success rate and instruction tokens or max subtree depth with statistical significance, the trends for other statistics may differ among the methods.
@@
Appendix K Finetuning May Generally Outperform Prompting in Compositional Tasks
In this section, we examine the capability of finetuned LLMs and prompted LLMs for compositional generalization in natural language reasoning tasks, rather than agentic tasks. We employ Last Letter Concatenation (Wei et al., 2022; Zhou et al., 2023a) as a benchmark, where if LLMs take “hans, structure” as a question, LLMs are asked to answer it as The last letter of “hans” is “s”. The last letter of “structure” is “e”. Considering “s”, “e” leads to “se”. So, “hans, structure” outputs “se”. This reasoning problem exhibits a compositional nature as we increase the number of words provided as a question. For finetuned LLMs, we adopt 11B-parameter Flan-T5-XXL (Chung et al., 2022; Raffel et al., 2020) as a base model, and prepare 2000 examples per $n$-letter for a training dataset. For prompted LLMs, we employ Flan-PaLM-8B, 62B, and 540B (Chung et al., 2022; Chowdhery et al., 2022), and prepare randomized 4-shot CoT prompts from the same training dataset. We randomly use one of them during inference. Furthermore, we use up to 8 letters as a training dataset and few-shot exemplars and also use 9-12 letters as a hold-out test dataset to investigate the compositional generalization in natural language reasoning tasks.
Table 17 shows that finetuned LLM (Flan-T5-XXL) almost always outperforms prompted LLM (Flan-U-PaLM-540B) on average (76% v.s. 44% in an exact match) and also in both in-distribution (2-8 letters) and out-of-distribution (9-12 letters) settings. This implies that finetuned LLMs may generally outperform prompted LLMs in compositional tasks, which can also lead to the better performance of finetuned LMAs than prompted LMAs in sequential web automation tasks, as shown in the main text.
# of Letters
Models 2 3 4 5 6 7 8 9 10 11 12 Ave.
Flan-PaLM-8B (Prompted) 0.10 0.01 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.01
Flan-PaLM-62B (Prompted) 0.48 0.16 0.05 0.03 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.07
Flan-U-PaLM-540B (Prompted) 1.00 1.00 0.75 0.76 0.57 0.41 0.16 0.08 0.06 0.00 0.01 0.44
Flan-T5-XXL (11B; Finetuned) 1.00 0.98 0.99 0.98 0.98 0.96 0.96 0.83 0.47 0.20 0.05 0.76
Table 17: The performance of finetuning (Flan-T5-XXL (Chung et al., 2022; Raffel et al., 2020), 11B parameters) and prompting (Flan-PaLM-8B/62B/540B (Chung et al., 2022; Chowdhery et al., 2022)) in Last Letter Concatenation (Wei et al., 2022; Zhou et al., 2023a). We measure the performance with an exact match. We use up to 8 letters as a training dataset and few-shot exemplars and also use 9-12 letters as a hold-out test dataset to investigate the compositional generalization in natural language reasoning tasks. The results show that finetuned LLM (Flan-T5-XXL) almost always outperforms prompted LLM (Flan-U-PaLM-540B) in both in-distribution (2-8 letters) and out-of-distribution (9-12 letters) settings.